OpenStudy (anonymous):

What is the quadratic function that is created with roots -10 and -6 and a vertex at (-8, -8)?

OpenStudy (anonymous):

(x+10) and (x+6) are the zeros.

OpenStudy (anonymous):

it should = 16x

OpenStudy (anonymous):

Factors: y=(x+10)(x+6) FOIL to get y=x^2+16x+60 Complete the Square: y=(x^2+16x+8)+60-8

OpenStudy (anonymous):

but isn't that giving me a vertex of -8,-4